OpenStudy (anonymous):

If (x + 1)(x - 5) = -9, then which of the following statements is true? x + 1 = 0 or x - 5 = 0 x + 1 = -9 or x - 5 = -9 x - 2 = 0

OpenStudy (anonymous):

im thinking its the second one

jimthompson5910 (jim_thompson5910):

(x + 1)(x - 5) = -9 x^2 - 5x + x - 5 = -9 x^2 - 5x + x - 5 + 9 = 0 x^2 - 4x + 4 = 0 (x - 2)(x - 2) = 0 x - 2 = 0 or x - 2 = 0 x - 2 = 0 So it's the last choice

myininaya (myininaya):

Yeah you can only do a*b=0 implies a=0 or b=0 or both=0

myininaya (myininaya):

You have to have that product =0

OpenStudy (anonymous):

oh i see

OpenStudy (anonymous):

thank u